OpenStudy (anonymous):

What is the seventh term in the addition pattern that begins 17, 32, 47, 62? A. 107 B. 117 C. 122 D. 137

OpenStudy (anonymous):

@thomaster @TrevorBaker @LifeIsADangerousGame @coolman500 @kropot72 @leonardo0430

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (anonymous):

B because look at the first number brother

OpenStudy (anonymous):

sister you made me a 40% last time...

OpenStudy (anonymous):

N o i didnt i only made you a 0 in gum dummy

OpenStudy (anonymous):

That doesn't mean anything mathgum

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

whats the pattern?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@dan815 @Jaynator495 @Daniellelovee

OpenStudy (anonymous):

\[17+15=32+15=?\]

OpenStudy (anonymous):

47

OpenStudy (anonymous):

its a patteren u add 15 to each number till u get to the seventh number

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

107 is the seventh number so 107 is ur answer

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

np